OpenStudy (anonymous):

If h(x) = x − 1 and j(x) = −4x, find h[j(5)]. −19 −21 −16 −20

OpenStudy (anonymous):

-21

OpenStudy (anonymous):

j(5)= -4(5) = -20 then h(-20 )= -20-1 = -21